# Projective simplicity and Bergman’s property for unit groups of continuous rings

### v1

@@ 2. Preliminaries
A commutator in $G$ is any element of the form $g,h]\mathrel{\mathop{:}}=g^{-1}h^{-1}gh$ where $g,h\in G$. For subgroups $H_{0},H_{1}\leq G$, we define
@@ 4. A general criterion
    There is a sequence $(e_{n})_{n\in\operatorname{\mathbb{N}}}$ of pairwise orthogonal elements of $E$ such that, for every sequence $h_{n}\in G(e_{n})$ $(n\in\operatorname{\mathbb{N}})$, there exists $g\in G$ such that
@@
    For every $e\in E$, there exist a finite subset $F\subseteq G(e)$, $m\in\operatorname{\mathbb{N}}$ and $f\in E$ such that $f\leq e$ and
@@
Claim 1. There exist $m_{1},n_{1}\in\operatorname{\mathbb{N}}$ and $f\in E$ such that $f\leq e$ and $G(f)\subseteq W_{n_{1}}^{m_{1}}$.
Proof of Claim 1. By (B), there exist a finite set $F_{1}\subseteq G(e)$, $m\in\operatorname{\mathbb{N}}$ and $f\in E$ such that $f\leq e$ and $G(f)\subseteq\operatorname{Cl}_{G(e)}(F_{1})^{m}$. As $F_{1}$ is finite, we find $k_{1}\in\operatorname{\mathbb{N}}$ with $F_{1}\subseteq W_{k_{1}}$. Let $n_{1}\mathrel{\mathop{:}}=\max\{n_{0},k_{1}\}$ and $m_{1}\mathrel{\mathop{:}}=3m$. We will show that $G(f)\subseteq W_{n_{1}}^{m_{1}}$. To this end, let $h\in G(f)$. Then there exist $g_{1},\ldots,g_{m}\in F_{1}$ and $t_{1},\ldots,t_{m}\in G(e)$ such that $h=t_{1}g_{1}t_{1}^{-1}\cdots t_{m}g_{m}t_{m}^{-1}$. Since $W_{n_{0}}$ is full for $e$, for each $i\in\{1,\ldots,m\}$ there exists $s_{i}\in W_{n_{0}}$ such that $es_{i}=s_{i}e=t_{i}e$, or equivalently, $s_{i}^{-1}e=es_{i}^{-1}=t_{i}^{-1}e$. Now,
@@
Therefore, as desired,
@@ 5. Uncountable strong cofinality
Lemma 5.5 (cf. [7, Proposition 5] [^7] In [7, Proposition 5], this is stated under the negligible hypothesis of $R$ being finitely generated.).
@@
This shows that $\operatorname{GL}_{n+1}(R)\subseteq H_{1}GH_{2}H_{1}H_{2}$, which entails the desired equality. ∎
@@
$H_{1}GH_{2}H_{1}H_{2}\,\subseteq\,\left(G\cup\left\{c,c^{-1}\right\}\right)^{19}.$
@@
$H_{1}GH_{2}H_{1}H_{2}\,\subseteq\,c^{-1}Gc^{2}Gc^{-1}GcGc^{-2}Gc^{2}Gc^{-2}Gc\,\subseteq\,\left(G\cup\left\{c,c^{-1}\right\}\right)^{19}.\qed$
@@
Then $\operatorname{GL}_{n+1}(R)=\left(G\cup\left\{c,c^{-1}\right\}\right)^{19}$.
@@
$\Gamma_{R}(f)\,=\,\left(\Gamma_{R}(e)\cup\left\{g,g^{-1}\right\}\right)^{19}.$
@@
$\operatorname{GL}(S)\,$ $=\,\varphi(\operatorname{GL}_{3}(T))\,=\,\varphi\!\left(\left(G\cup\left\{c,c^{-1}\right\}\right)^{19}\right)$
$=\,\left(\varphi(G)\cup\left\{\varphi(c),\varphi(c)^{-1}\right\}\right)^{19}\,=\,\left(\Gamma_{S}(e)\cup\left\{\varphi(c),\varphi(c)^{-1}\right\}\right)^{19}.$
@@
$\Gamma_{R}(f)\,$ $=\,\iota(\operatorname{GL}(S))\,=\,\iota\!\left(\left(\Gamma_{S}(e)\cup\left\{\varphi(c),\varphi(c)^{-1}\right\}\right)^{19}\right)$
$=\,\left(\iota(\Gamma_{S}(e))\cup\left\{\iota(\varphi(c)),\iota(\varphi(c))^{-1}\right\}\right)^{19}\,=\,\left(\Gamma_{R}(e)\cup\left\{g,g^{-1}\right\}\right)^{19}$
@@
$\Gamma_{R}(e_{i-1})\,=\,\left(\Gamma_{R}(e_{i})\cup\left\{g_{i},g_{i}^{-1}\right\}\right)^{19}.$
Hence, considering $\ell\mathrel{\mathop{:}}=19^{n}$ and $F\mathrel{\mathop{:}}=\left\{g_{1},\ldots,g_{n},g_{1}^{-1},\ldots,g_{n}^{-1}\right\}$, we infer that
$\operatorname{GL}(R)\,$ $=\,\Gamma_{R}(e_{0})\,=\,\left(\Gamma_{R}(e_{1})\cup\left\{g_{1},g_{1}^{-1}\right\}\right)^{19}\,\subseteq\,\left(\Gamma_{R}(e_{2})\cup\left\{g_{1},g_{2},g_{1}^{-1},g_{2}^{-1}\right\}\right)^{19^{2}}$

### v2

@@ 2. Preliminaries
A subgroup $H_{0}\leq G$ is said to be normalized by a subgroup $H_{1}\leq G$ if $gH_{0}g^{-1}=H_{0}$ for every $g\in H_{1}$. A commutator in $G$ is any element of the form $g,h]\mathrel{\mathop{:}}=g^{-1}h^{-1}gh$ where $g,h\in G$. For subgroups $H_{0},H_{1}\leq G$, we define
@@ 4. A general criterion
    There exists $(e_{n})_{n\in\operatorname{\mathbb{N}}}\in E^{\operatorname{\mathbb{N}}}$ such that, for every sequence $h_{n}\in G(e_{n})$ $(n\in\operatorname{\mathbb{N}})$, there exists $g\in G$ such that
@@
    For every $e\in E$, there exist a finite subset $F\subseteq G(e)$, $m\in\operatorname{\mathbb{N}}$ and $f\in E$ with
@@
Claim 1. There exist $m_{1},n_{1}\in\operatorname{\mathbb{N}}$ and $f\in E$ such that $G(f)\subseteq W_{n_{1}}^{m_{1}}$.
Proof of Claim 1. By (B), there exist a finite set $F_{1}\subseteq G(e)$, $m\in\operatorname{\mathbb{N}}$ and $f\in E$ such that $G(f)\subseteq\operatorname{Cl}_{G(e)}(F_{1})^{m}$. As $F_{1}$ is finite, we find $k_{1}\in\operatorname{\mathbb{N}}$ such that $F_{1}\subseteq W_{k_{1}}$. Let $n_{1}\mathrel{\mathop{:}}=\max\{n_{0},k_{1}\}$ and $m_{1}\mathrel{\mathop{:}}=3m$. We will show that $G(f)\subseteq W_{n_{1}}^{m_{1}}$. To this end, let $h\in G(f)$. Then there exist $g_{1},\ldots,g_{m}\in F_{1}$ and $t_{1},\ldots,t_{m}\in G(e)$ such that $h=t_{1}g_{1}t_{1}^{-1}\cdots t_{m}g_{m}t_{m}^{-1}$. Since $W_{n_{0}}$ is full for $e$, for each $i\in\{1,\ldots,m\}$ there exists $s_{i}\in W_{n_{0}}$ such that $es_{i}=s_{i}e=t_{i}e$, or equivalently, $s_{i}^{-1}e=es_{i}^{-1}=t_{i}^{-1}e$. Now,
@@
Therefore, as $h\in G(f)\subseteq\operatorname{Cl}_{G(e)}(F_{1})^{m}\subseteq G(e)$,
@@ 5. Uncountable strong cofinality
Lemma 5.5 (cf. [7, Proposition 5] [^7] The statement of [7, Proposition 5] contains the negligible hypothesis of $R$ being finitely generated.).
@@
This shows that $\operatorname{GL}_{n+1}(R)=H_{1}GH_{2}H_{1}H_{2}$. Since
$\begin{pmatrix}a&0\\ 0&1\end{pmatrix}\begin{pmatrix}1&x\\ 0&1\end{pmatrix}{\begin{pmatrix}a&0\\ 0&1\end{pmatrix}}^{-1}\!\,=\,\begin{pmatrix}a&0\\ 0&1\end{pmatrix}\begin{pmatrix}1&x\\ 0&1\end{pmatrix}\begin{pmatrix}a^{-1}&0\\ 0&1\end{pmatrix}\,=\,\begin{pmatrix}1&ax\\ 0&1\end{pmatrix}\,\in\,H_{1}$
for all $a\in\operatorname{GL}_{n}(R)$ and $x\in R^{n\times 1}$, we know that $H_{1}$ is normalized by $G$, thus
$\operatorname{GL}_{n+1}(R)\,=\,H_{1}GH_{2}H_{1}H_{2}\,=\,GH_{1}H_{2}H_{1}H_{2}.\qed$
@@
$GH_{1}H_{2}H_{1}H_{2}\,\subseteq\,\left(G\cup\left\{c,c^{-1}\right\}\right)^{16}.$
@@
$GH_{1}H_{2}H_{1}H_{2}\,\subseteq\,Gc^{-1}Gc^{2}Gc^{-2}Gc^{2}Gc^{-2}Gc\,\subseteq\,\left(G\cup\left\{c,c^{-1}\right\}\right)^{16}.\qed$
@@
Then $\operatorname{GL}_{n+1}(R)=\left(G\cup\left\{c,c^{-1}\right\}\right)^{16}$.
@@
$\Gamma_{R}(f)\,=\,\left(\Gamma_{R}(e)\cup\left\{g,g^{-1}\right\}\right)^{16}.$
@@
$\operatorname{GL}(S)\,$ $=\,\varphi(\operatorname{GL}_{3}(T))\,=\,\varphi\!\left(\left(G\cup\left\{c,c^{-1}\right\}\right)^{16}\right)$
$=\,\left(\varphi(G)\cup\left\{\varphi(c),\varphi(c)^{-1}\right\}\right)^{16}\,=\,\left(\Gamma_{S}(e)\cup\left\{\varphi(c),\varphi(c)^{-1}\right\}\right)^{16}.$
@@
$\Gamma_{R}(f)\,$ $=\,\iota(\operatorname{GL}(S))\,=\,\iota\!\left(\left(\Gamma_{S}(e)\cup\left\{\varphi(c),\varphi(c)^{-1}\right\}\right)^{16}\right)$
$=\,\left(\iota(\Gamma_{S}(e))\cup\left\{\iota(\varphi(c)),\iota(\varphi(c))^{-1}\right\}\right)^{16}\,=\,\left(\Gamma_{R}(e)\cup\left\{g,g^{-1}\right\}\right)^{16}$
@@
$\Gamma_{R}(e_{i-1})\,=\,\left(\Gamma_{R}(e_{i})\cup\left\{g_{i},g_{i}^{-1}\right\}\right)^{16}.$
Hence, considering $\ell\mathrel{\mathop{:}}=16^{n}$ and $F\mathrel{\mathop{:}}=\left\{g_{1},\ldots,g_{n},g_{1}^{-1},\ldots,g_{n}^{-1}\right\}$, we infer that
$\operatorname{GL}(R)\,$ $=\,\Gamma_{R}(e_{0})\,=\,\left(\Gamma_{R}(e_{1})\cup\left\{g_{1},g_{1}^{-1}\right\}\right)^{16}\,\subseteq\,\left(\Gamma_{R}(e_{2})\cup\left\{g_{1},g_{2},g_{1}^{-1},g_{2}^{-1}\right\}\right)^{16^{2}}$